Fifteen dogs are three seventeen of the number of dogs certain living area. How many dogs in total are there in the living area?


1
Expert's answer
2021-05-24T16:14:38-0400

If "\frac{3}{17}=15dogs"

"\frac{17}{17}=?\:"


"Total\:number\:of\:dogs=\frac{17}{17}\:\times 15\times \frac{17}{3}=85"

Therefore, there are 85 dogs in total, in the living area.